# Momentum coupling of classical atoms

### v2

@@ 1 Introduction
The derivation of van der Waals’ equation of state was first realized in one dimension. Previously, Kac [6] had shown how to calculate the grand partition function for a one-dimensional gas with a potential constructed from a hard core and attraction between particles at $x_{i}$ and $x_{j}$ of the form $-\alpha e^{-\gamma|x_{i}-x_{j}|}$. In this model there is no phase transition, however, by replacing $\alpha$ with $\alpha_{0}\gamma$ and sending $\gamma$ to zero after the thermodynamic limit, van der Waals’ equation of state as well as Maxwell’s equal-area rule could be reproduced [7]. The method applied was dimension-specific, the same result in higher dimensions required new ideas, partly originating from van Kempen [8], and was obtained by Lebowitz and Penrose [10].
@@
In this paper we propose a new approach to the problem of phase transitions in systems of classical particles in the continuum. It can be called the dual-space cluster expansion. The interaction, which we restrict to translation-invariant pair potentials, acts by correlating the motions of the particles. For interactions that decay sufficiently rapidly with the distance the correlation is weak at low density. By raising the density a qualitative change in the correlated motion of atoms can take place. The position and the momentum of classical particles are independent, the partition function factors accordingly, and any phase transition can arise only from the position-dependent configurational partition function, an integral of the $N$-particle Boltzmann factor over the different configurations of $N$ motionless particles in a bounded volume. By Fourier-expanding the two-particle Boltzmann factors and integrating with respect to the positions one can introduce a new set of variables, the momenta of the relative displacements of particle pairs. For $N$ position vectors there are $N(N-1)/2$ momenta, so they are subject to constraints, the physical meaning of which is to ensure that the new variables still describe a system of immobile particles. Analysis of the constraints reveals that they are of two types, graphical and numerical. A set of momenta can be represented on a graph whose $N$ labeled vertices indicate the particles and whose edges correspond to the nonzero momenta. A properly directed momentum vector is assigned to every edge, and at each vertex the sum of the incident momenta – the total momentum of the corresponding particle – must be zero. A graph on which this condition can be satisfied is called valid. We will see that the valid graphs are cyclic, that is, each edge is in at least one cycle. The graph is decomposed into maximal connected subgraphs, called clusters. Our gain from the Fourier expansion is precisely this decomposition, allowing us to derive a formula for the density of particles in finite clusters in the infinite system. The question then is whether, at low density, there are only finite clusters, and whether, as the density increases, infinite clusters appear. Such a transition is reminiscent of percolation in lattice models, but there could be more than one infinite cluster, and the singularity at the transition point would be stronger.
@@
We will show that indeed, at low density there are only finite clusters. The appearance of infinite clusters depends on the convergence of a fugacity expansion at the convergence radius. The difficulty to prove it is partly combinatorial, we would need a good knowledge of the statistics of connected cyclic graphs on $n$ labeled vertices. An additional difficulty is that the coefficients of the fugacity expansion are obtained as a sum over these graphs, and interaction enters the sum in a complex way. The best we can do is to preserve only the $(n-1)!/2$ cycles of length $n$. With this approximation, for a repulsive pair potential decaying as a Gaussian at large distances and logarithmically diverging at the origin, for dimensions two and beyond we find that infinite clusters appear at high densities. Furthermore, without any approximation we can still prove that condensed phases contain infinite clusters. The vapor-liquid transition can be associated with the appearance of infinitely many infinite but submacroscopic clusters; on the other hand, crystallization is expected to be accompanied by the formation of one or more macroscopic clusters.
@@ 2 Recurrence formula for the partition function
exists and is continuous. We will need more, $\hat{v}\in L^{1}({\mathbb{R}}^{d})$, so both $v$ and $\hat{v}$ are continuous. $-v$ is the central object in the cluster expansion according to Mayer [12]. Here its Fourier transform will play the key role.
The complete canonical partition function for $N$ identical particles on the cube $\Lambda=[-L/2,L/2]^{d}$ with periodic boundary conditions is
$Q_{N,L}=\frac{1}{\lambda_{\beta}^{dN}N!}\int_{\Lambda^{N}}{\,\rm d}x_{1}\dots\,\rm dx_{N}\prod_{1\leq i<j\leq N}E_{L}(x_{j}-x_{i}).$
@@
Because $Z_{k}=0$ for $k=1,\dots,N$ must be satisfied, not all the $N(N-1)/2$ variables $z^{j}_{i}$ can be chosen freely. Those taking zero do not influence the value of $Z_{k}$’s, so one must focus on the set of nonzero $z^{j}_{i}$. Given $\{z^{j}_{i}\in{\mathbb{Z}}^{d}|1\leq i<j\leq N\}$, consider therefore a graph $G_{\{z^{j}_{i}\}}$ of $N$ vertices labeled from 1 to $N$ and an edge between $i$ and $j$ if $z^{j}_{i}\neq 0$. $G_{\{z^{j}_{i}\}}$ divides into clusters (components, in graph theory). Vertex $k$ is a single-point cluster if $z^{k}_{i}=0$ and $z^{j}_{k}=0$ for all $i<k<j$. We call a nontrivial cluster valid if $Z_{k}=0$ can be solved for each $k$ in its vertex set with all the edge variables taking a nonzero value. A graph is valid if all of its nontrivial clusters are valid. Below we use the standard notation $V(G)$ and $E(G)$ for the sets of vertices and edges, respectively, of a graph $G$, and $|V(G)|$ and $|E(G)|$ for the number of elements in these sets.
@@
where $x_{0,1}=x_{n,n+1}=x_{n,1}$. For clarity, we have written four equations instead of just two independent ones. We see that whatever be the choice of, say, $x_{1,2}$, the other variables must take the same value with plus or minus sign. So there is at most a single free variable. To be definite, let $l_{1}=\min_{j}l_{j}$. Let us take an arbitrary nonzero vector $v$ and set $x_{1,2}=v$. $Z_{l_{1}}=0$ is equation (1) for $i=1$, it is solved with $x_{n,1}=-v$. We must check that going around the cycle there is no ”frustration”, all the equations can be solved. It is helpful to imagine an arrow on every edge, pointing towards the larger-labeled vertex. Vertex $i$ is called a source if $l_{i}<l_{i-1},l_{i+1}$ (two outgoing arrows), a sink if $l_{i}>l_{i-1},l_{i+1}$ (two incoming arrows), and a transmitting point if $l_{i-1}<l_{i}<l_{i+1}$ or $l_{i-1}>l_{i}>l_{i+1}$ (one arrow in, one out). The problem is soluble because the number of sinks equals the number of sources. $i=1$ is a source; by passing it, $-v$ changes to $v$ while solving equation (1), stays $v$ by passing transmitting points until the first sink and solves equation (2), passing the sink it changes to $-v$ while solving equation (3), stays $-v$ by passing transmitting points and solves equation (4), and so on.
@@
the other edges keep their individual value. Clearly, $Z_{k}=0$ still holds for all $k\in\{i_{1},\dots,i_{m}\}\cup\{l_{1},\dots,l_{n}\}$, and all the edge variables are nonzero. The procedure can be continued by merging further cycles, and in a finite number of steps one can obtain any given connected finite cyclic graph.
In the opposite direction, suppose that $C$ is a valid cluster, and let a nonzero vector be assigned to each edge so that $Z_{k}=0$ for every $k\in V(C)$. Being valid, $C$ is without endpoints: if $k$ were an endpoint, it would have a single neighbor $j$, so $Z_{k}=z^{k}_{j}=0$ or $Z_{k}=z^{j}_{k}=0$ would be the only solution. Thus, $C$ contains cycles, and we can proceed by successive demerging. We choose a cycle $g$ in $C$ and prepare an image $g^{\prime}$ of $g$ outside $C$, meaning that $g^{\prime}$ has the same labeled vertices and the same edges as $g$. Next, we choose an edge $e$ of $g$ and, if it carries the vector $x$, we assign $-x$ to the image $e^{\prime}$ of $e$. This uniquely determines the variables on the other edges of $g^{\prime}$ in such a way that the constraints are satisfied. Now we merge $g^{\prime}$ with $C$. As a result, the new vector on $e$ is zero, and $x$ or $-x$ must be added to the vectors on the other edges of $g$. By dropping all the edges from the merger whose new vector is zero and by dropping also the vertices that become isolated, we obtain a new graph, not necessarily connected, having at least one edge less than $C$, and still being valid. In a finite number of steps we can eliminate all the cycles of $C$. If the rest is nonempty, it cannot be valid, which is a contradiction. So $C$ is cyclic. $\blacksquare$
The two products in Eq. (2) split into factors according to the clusters. We rearrange the sum constituting $Q_{N,L}$ by summing first over the clusters containing vertex 1. Due to permutation invariance, for $n$-point clusters it suffices to consider the vertex set $\{1,2,\dots,n\}$ and to take into account all the possible partners of 1 in a cluster of $n$ points by a factor (${N-1\choose n-1}$. Let $\Sigma_{n}$ denote the family of valid $n$-point clusters of vertex set $\{1,2,\dots,n\}$. The only element of $\Sigma_{1}$ is $\{1\}$, $\Sigma_{2}$ is empty. For $n\geq 3$ the number of edges are between $n$ and $n(n-1)/2$. The clusters with $n$ edges are cycles, their number is $(n-1)!/2$, and there is a single cluster of $n(n-1)/2$ edges, the complete $n$-graph. Separating in $Q_{N,L}$ the contribution of clusters of order $n$ containing 1, there remains $Q_{N-n,L}$. With $\rho=N/L^{d}$,
$Q_{N,L}$ $=$ $\frac{L^{dN}}{\lambda_{\beta}^{dN}N!}\sum_{n=1}^{N}{N-1\choose n-1}\hat{E}_{L}(0)^{n(N-n)}\sum_{C\in\Sigma_{n}}\hat{E}_{L}(0)^{{n\choose 2}-|E(C)|}\sum_{\{z^{j}_{i}\in{\mathbb{Z}}^{d}\setminus\{0\}|(ij)\in E(C)\}}\left(\prod_{(ij)\in E(C)}\hat{E}_{L}(z^{j}_{i})\right)\prod_{i=1}^{n}\delta_{Z_{i},0}$ (3)
@@
$=$ $\frac{1}{\rho\lambda_{\beta}^{d}}\sum_{n=1}^{N}\hat{E}_{L}(0)^{n(N-n)}q_{n}^{L}Q_{N-n,L},$
where $q^{L}_{1}=1$, $q^{L}_{2}=0$ and, for $n\geq 3$,
@@
Observe that $q^{L}_{n}$ is independent of $N$ and can be negative. Equation (3) is the recurrence relation announced in the title of the section.
@@ 3 Thermodynamic limit
The equality must hold when $N,L\to\infty$ while $N/L^{d}=\rho$ is fixed. For a positive integer $M<N$ we cut the sum in two, for $n\leq M$ and $M+1\leq n\leq N$. Taking the thermodynamic limit,
@@
$\frac{{Q}_{N-1,L}}{{Q}_{N,L}}=\exp\left\{\beta\frac{f_{N,L}(\beta)-f_{N-1,L}(\beta)}{N/L^{d}-(N-1)/L^{d}}\right\}\to\exp\left\{\beta\frac{\partial{f}(\rho,\beta)}{\partial\rho}\right\},$
@@
$f(\rho,\beta)$ is the free energy density, $\partial f(\rho,\beta)/\partial\rho$ is the chemical potential, and $\exp\{\beta\partial f(\rho,\beta)/\partial\rho\}$ is the fugacity. Furthermore,
@@
Let us see, how infinite clusters can emerge at high density. $A_{\rho,\beta}(\mu)$ must be finite for the equation (10) to be solved. The sum is absolutely convergent if $\mu<\bar{\mu}(\rho,\beta)$, where $e^{\beta\bar{\mu}(\rho,\beta)}$ is the convergence radius of the power series $\sum_{n}e^{-\rho\hat{v}(0)}q_{n}z^{n}$. Stability of the interaction implies that $|q_{n}|$ cannot increase faster than exponentially with $n$, therefore the convergence radius is nonzero, $\bar{\mu}(\rho,\beta)>-\infty$. Fixing any $\beta<\infty$, for $\rho$ sufficiently small, Eq. (10) can be solved for $\mu$. If $\mu(\rho,\beta)$ denotes the solution, then $A_{\rho,\beta}(\mu)=0$ yields $\mu(0,\beta)=-\infty$ and, due to the convexity of $f(\rho,\beta)$ as a function of $\rho$ [14], $\mu(\rho,\beta)=\partial f(\rho,\beta)/\partial\rho$ strictly increases with $\rho$ in an interval $(0,\rho_{c}(\beta))$, where $\rho_{c}(\beta)\leq\infty$. Taking $\mu$ as an independent variable, $A_{\rho,\beta}(\mu)$ strictly increases with it in the interval $(-\infty,\bar{\mu}(\rho,\beta))$, and tends to $A_{\rho,\beta}(\bar{\mu}(\rho,\beta))\leq\infty$. If $A_{\rho,\beta}(\bar{\mu}(\rho,\beta))=\infty$ then $\rho_{c}(\beta)=\infty$. If $A_{\rho,\beta}(\bar{\mu}(\rho,\beta))$ is finite, $\rho_{c}(\beta)$ is also finite.
Now whether $A_{\rho,\beta}(\bar{\mu}(\rho,\beta))$ is finite or not depends on the asymptotic behavior of $q_{n}$. If
@@
$A_{\rho,\beta}(\mu)=e^{\beta\mu-\rho\hat{v}(0)}+\sum_{n=3}^{\infty}\alpha_{n}e^{-n\beta\delta_{n}(\beta)}e^{n[\beta\mu-\beta\epsilon(\beta)-\rho\hat{v}(0)]}.$ (12)
Because $n\delta_{n}(\beta)=o(n)$, $\beta\mu-\beta\epsilon(\beta)-\rho\hat{v}(0)=0$ determines
@@
independent of $\rho$, even though $\bar{\mu}(\rho,\beta)$ depends on it. We see that the fate of infinite clusters depends on the correction $e^{o(n)}$ to the exponential decay or growth of $|q_{n}|$; the whole difficulty of proving the transition lies in this fact. In this respect, the appearance of infinite clusters is analogous to the appearance of infinite permutation cycles in systems of interacting bosons [18]. It cannot be excluded that, for some interactions, $|q_{n}|\sim\exp\{-c_{1}n^{1+c_{2}}\}$ with $c_{1},\,c_{2}>0$. Then, the radius of convergence of the series $A_{\rho,\beta}(\mu)$ is infinite, there are only finite clusters and no classical phase transitions at all densities. This may be the case for long-range repulsive forces, which – as previously noted [18] – can also prevent the occurrence of infinite permutation cycles, and thus of Bose-Einstein condensation.

### v3

@@ 1 Introduction
The derivation of van der Waals’ equation of state was first realized in one dimension. Previously, Kac [6] had shown how to calculate the grand partition function for a one-dimensional gas with a potential constructed from a hard core and attraction between particles at $x_{i}$ and $x_{j}$ of the form $\alpha e^{-\gamma|x_{i}-x_{j}|}$. In this model there is no phase transition, however, by replacing $\alpha$ with $\alpha_{0}\gamma$ and sending $\gamma$ to zero after the thermodynamic limit, van der Waals’ equation of state as well as Maxwell’s equal-area rule could be reproduced [7]. The method applied was dimension-specific, the same result in higher dimensions required new ideas, partly originating from van Kempen [8], and was obtained by Lebowitz and Penrose [10].
@@
In this paper we propose a new approach to the problem of phase transitions in systems of classical particles in the continuum. It can be called the dual-space cluster expansion. The interaction, which we restrict to translation-invariant pair potentials, acts by correlating the motions of the particles. For interactions that decay sufficiently rapidly with the distance the correlation is weak at low density. By raising the density a qualitative change in the correlated motion of atoms can take place. The position and the momentum of classical particles are independent, the partition function factors accordingly, and any phase transition can arise only from the position-dependent configurational partition function, an integral of the $N$-particle Boltzmann factor over the different configurations of $N$ motionless particles in a bounded volume. By Fourier-expanding the two-particle Boltzmann factors and integrating with respect to the positions one can introduce a new set of variables, the momenta of the relative displacements of particle pairs. For $N$ position vectors there are N(N1)/2N(N-1)/2 momenta, so they are subject to constraints, the physical meaning of which is to ensure that the new variables still describe a system of immobile particles. Analysis of the constraints reveals that they are of two types, graphical and numerical. A set of momenta can be represented on a graph whose $N$ labeled vertices indicate the particles and whose edges correspond to the nonzero momenta. A properly directed momentum vector is assigned to every edge, and at each vertex the sum of the incident momenta – the total momentum of the corresponding particle – must be zero. A graph on which this condition can be satisfied is called valid. We will see that the valid graphs are cyclic, that is, each edge is in at least one cycle. The graph is decomposed into maximal connected subgraphs, called clusters. Our gain from the Fourier expansion is precisely this decomposition, allowing us to derive a formula for the density of particles in finite clusters in the infinite system. The question then is whether, at low density, there are only finite clusters, and whether, as the density increases, infinite clusters appear. Such a transition is reminiscent of percolation in lattice models, but there could be more than one infinite cluster, and the singularity at the transition point would be stronger.
@@
We will show that indeed, at low density there are only finite clusters. The appearance of infinite clusters depends on the convergence of an activity expansion at the convergence radius. The difficulty to prove it is partly combinatorial, we would need a good knowledge of the statistics of connected cyclic graphs on $n$ labeled vertices. An additional difficulty is that the coefficients of the activity expansion are obtained as a sum over these graphs, and interaction enters the sum in a complex way. The best we can do is to preserve only the (n1)!/2(n-1)!/2 cycles of length $n$. With this approximation, for a repulsive pair potential decaying as a Gaussian at large distances and logarithmically diverging at the origin, for dimensions two and beyond we find that infinite clusters appear at high densities. Furthermore, without any approximation we can still prove that condensed phases contain infinite clusters. The vapor-liquid transition can be associated with the appearance of infinitely many infinite but submacroscopic clusters; on the other hand, crystallization is expected to be accompanied by the formation of one or more macroscopic clusters.
@@ 2 Recurrence formula for the partition function
exists and is continuous. We will need more, $\hat{v}\in L^{1}({\mathbb{R}}^{d})$, so both $v$ and $\hat{v}$ are continuous. v-v is the central object in the cluster expansion according to Mayer [12]. Here its Fourier transform will play the key role.
The complete canonical partition function for $N$ identical particles on the hypercube $\Lambda=[-L/2,L/2]^{d}$ with periodic boundary conditions is
$Q_{N,L}=\frac{1}{\lambda_{\beta}^{dN}N!}\int_{\Lambda^{N}}{\,\rm d}x_{1}\dots\,\rm dx_{N}\prod_{1\leq i<j\leq N}E_{L}(x_{j}-x_{i}).$ (1)
@@
Because $Z_{k}=0$ for $k=1,\dots,N$ must be satisfied, not all the N(N1)/2N(N-1)/2 variables $z^{j}_{i}$ can be chosen freely. Those taking zero do not influence the value of $Z_{k}$’s, so one must focus on the set of nonzero $z^{j}_{i}$. Given $\{z^{j}_{i}\in{\mathbb{Z}}^{d}|1\leq i<j\leq N\}$, consider therefore a graph $G_{\{z^{j}_{i}\}}$ of $N$ vertices labeled from 1 to $N$ and an edge between $i$ and $j$ if $z^{j}_{i}\neq 0$. $G_{\{z^{j}_{i}\}}$ divides into clusters (components, in graph theory). Vertex $k$ is a single-point cluster if $z^{k}_{i}=0$ and $z^{j}_{k}=0$ for all $i<k<j$. We call a nontrivial cluster valid if $Z_{k}=0$ can be solved for each $k$ in its vertex set with all the edge variables taking a nonzero value. A graph is valid if all of its nontrivial clusters are valid. Below we use the standard notation $V(G)$ and $E(G)$ for the sets of vertices and edges, respectively, of a graph $G$, and $|V(G)|$ and $|E(G)|$ for the number of elements in these sets.
@@
where $x_{0,1}=x_{n,n+1}=x_{n,1}$. For clarity, we have written four equations instead of just two independent ones. We see that whatever be the choice of, say, $x_{1,2}$, the other variables must take the same value with plus or minus sign. So there is at most a single free variable. To be definite, let $l_{1}=\min_{j}l_{j}$. Let us take an arbitrary nonzero vector $v$ and set $x_{1,2}=v$. $Z_{l_{1}}=0$ is equation (1) for $i=1$, it is solved with $x_{n,1}=-v$. We must check that going around the cycle there is no ”frustration”, all the equations can be solved. It is helpful to imagine an arrow on every edge, pointing towards the larger-labeled vertex. Vertex $i$ is called a source if $l_{i}<l_{i-1},l_{i+1}$ (two outgoing arrows), a sink if $l_{i}>l_{i-1},l_{i+1}$ (two incoming arrows), and a permeable point if $l_{i-1}<l_{i}<l_{i+1}$ or $l_{i-1}>l_{i}>l_{i+1}$ (one arrow in, one out). In order to satisfy the respective equation, on different sides of a source (equation (1)) or a sink (equation (3)), $v$ must appear with different signs, while on both sides of a permeable point (equations (2) and (4)) the sign must be the same; see Fig. 1(a). The problem is soluble because the number of sinks equals the number of sources. Specifically, $i=1$ is a source; by passing it, v-v changes to $v$ while solving equation (1), stays $v$ by passing permeable points until the first sink and solves equation (2), passing the sink it changes to v-v while solving equation (3), stays v-v by passing permeable points and solves equation (4), and so on.
(a)1245235v-v-$v$-www(b)12453v-v-$v$-www
Figure 1. Merging two cycles through a common edge.
@@
the other edges keep their individual value (Fig.1(b)). Clearly, $Z_{k}=0$ still holds for all $k\in\{i_{1},\dots,i_{m}\}\cup\{l_{1},\dots,l_{n}\}$, and all the edge variables are nonzero. The procedure can be continued by merging further cycles, and in a finite number of steps one can obtain any given connected finite cyclic graph.
In the opposite direction, suppose that $C$ is a valid cluster, and let a nonzero vector be assigned to each edge so that $Z_{k}=0$ for every $k\in V(C)$. Being valid, $C$ is without endpoints: if $k$ were an endpoint, it would have a single neighbor $j$, so $Z_{k}=z^{k}_{j}=0$ or $Z_{k}=z^{j}_{k}=0$ would be the only solution. Thus, $C$ contains cycles, and we can proceed by successive demerging. We choose a cycle $g$ in $C$ and prepare an image $g^{\prime}$ of $g$ outside $C$, meaning that $g^{\prime}$ has the same labeled vertices and the same edges as $g$. Next, we choose an edge $e$ of $g$ and, if it carries the vector $x$, we assign x-x to the image $e^{\prime}$ of $e$. This uniquely determines the variables on the other edges of $g^{\prime}$ in such a way that the constraints are satisfied. Now we merge $g^{\prime}$ with $C$. As a result, the new vector on $e$ is zero, and $x$ or x-x must be added to the vectors on the other edges of $g$. By dropping all the edges from the merger whose new vector is zero and by dropping also the vertices that become isolated, we obtain a new graph, not necessarily connected, having at least one edge less than $C$, and still being valid. In a finite number of steps we can eliminate all the cycles of $C$. If the rest is nonempty, it cannot be valid, which is a contradiction. So $C$ is cyclic. $\blacksquare$
The two products in Eq. (3) split into factors according to the clusters. We rearrange the sum constituting $Q_{N,L}$ by summing first over the clusters containing vertex 1. Due to permutation invariance, for $n$-point clusters it suffices to consider the vertex set $\{1,2,\dots,n\}$ and to take into account all the possible partners of 1 in a cluster of $n$ points by a factor $N-1\choose n-1$}. Let $\Sigma_{n}$ denote the family of valid $n$-point clusters of vertex set $\{1,2,\dots,n\}$. The only element of $\Sigma_{1}$ is $\{1\}$, $\Sigma_{2}$ is empty. For $n\geq 3$ the number of edges are between $n$ and n(n1)/2n(n-1)/2. The clusters with $n$ edges are cycles, their number is (n1)!/2(n-1)!/2, and there is a single cluster of n(n1)/2n(n-1)/2 edges, the complete $n$-graph. Separating in $Q_{N,L}$ the contribution of clusters of order $n$ containing 1, there remains $Q_{N-n,L}$. With $\rho=N/L^{d}$,
$Q_{N,L}$ $=$ $\frac{L^{dN}}{\lambda_{\beta}^{dN}N!}\sum_{n=1}^{N}{N-1\choose n-1}\hat{E}(0)^{n(N-n)}\sum_{C\in\Sigma_{n}}\hat{E}_{L}(0)^{{n\choose 2}-|E(C)|}\sum_{\{z^{j}_{i}\in{\mathbb{Z}}^{d}\setminus\{0\}|(ij)\in E(C)\}}\left(\prod_{(ij)\in E(C)}\hat{E}_{L}(z^{j}_{i})\right)\prod_{i=1}^{n}\delta_{Z_{i},0}$ (4)
@@
$=$ $\frac{1}{\rho\lambda_{\beta}^{d}}\sum_{n=1}^{N}\hat{E}(0)^{n(N-n)}q_{n}^{L}Q_{N-n,L},$
where $q^{L}_{1}=1$, $q^{L}_{2}=0$, and for $n\geq 3$,
@@
The factor $\hat{E}(0)^{n(N-n)}$ comes from $z^{k}_{j}=0$ when $j\leq n<k\leq N$. Observe that $q^{L}_{n}$ is independent of $N$ and can be negative. Equation (4) is the recurrence relation announced in the title of the section.
@@ 3 Thermodynamic limit
The equality must hold when $N,L\to\infty$ while $N/L^{d}=\rho$ is fixed. For a positive integer $M<N$ we cut the sum into two parts, for $n\leq M$ and $M+1\leq n\leq N$. Taking the thermodynamic limit,
@@
$\frac{{Q}_{N-1,L}}{{Q}_{N,L}}=\exp\left\{\beta\frac{f_{N,L}(\beta)-f_{N-1,L}(\beta)}{N/L^{d}-(N-1)/L^{d}}\right\}\to\exp\left\{\beta\frac{\partial{f}(\rho,\beta)}{\partial\rho}\right\},$ (9)
@@
$f(\rho,\beta)$ is the free energy density, $\partial f(\rho,\beta)/\partial\rho$ is the chemical potential, and $\exp\{\beta\partial f(\rho,\beta)/\partial\rho\}$ is the activity. Furthermore,
@@
For nonnegative pair potentials equations (1) and (9) yield the trivial bounds
$\rho\lambda_{\beta}^{d}\leq e^{\beta\partial f(\rho,\beta)/\partial\rho}\leq\rho\lambda_{\beta}^{d}\,e^{\beta\rho\int u{\,\rm d}x}.$
The lower bound requires $u\geq 0$, the upper bound is always true due to Jensen’s inequality. Note that $\int u{\,\rm d}x\geq 0$ holds for stable interactions, even if $u$ is partly negative.
Let us see, how infinite clusters can emerge at high density. $A_{\rho,\beta}(\mu)$ must be finite for the equation (12) to be solved. The sum is absolutely convergent if $\mu<\bar{\mu}(\rho,\beta)$, where $e^{\beta\bar{\mu}(\rho,\beta)}$ is the convergence radius of the power series $\sum_{n}e^{-n\rho\hat{v}(0)}q_{n}z^{n}$. Stability of the interaction implies that $|q_{n}|$ cannot increase faster than exponentially with $n$, therefore the convergence radius is nonzero, $\bar{\mu}(\rho,\beta)>$-$\infty$. Fixing any $\beta<\infty$, for $\rho$ sufficiently small, Eq. (12) can be solved for $\mu$. If $\mu(\rho,\beta)$ denotes the solution, then $A_{\rho,\beta}(\mu)=0$ yields $\mu(0,\beta)=$-$\infty$ and, due to the convexity of $f(\rho,\beta)$ as a function of $\rho$ [14], $\mu(\rho,\beta)=\partial f(\rho,\beta)/\partial\rho$ strictly increases with $\rho$ in an interval $(0,\rho_{c}(\beta))$, where $\rho_{c}(\beta)\leq\infty$. Taking $\mu$ as an independent variable, $A_{\rho,\beta}(\mu)$ strictly increases with it in the interval $(-\infty,\bar{\mu}(\rho,\beta))$, and tends to $A_{\rho,\beta}(\bar{\mu}(\rho,\beta))\leq\infty$. If $A_{\rho,\beta}(\bar{\mu}(\rho,\beta))=\infty$ then $\rho_{c}(\beta)=\infty$. If $A_{\rho,\beta}(\bar{\mu}(\rho,\beta))$ is finite, $\rho_{c}(\beta)$ is also finite. Infinite clusters only appear in this case, for $\rho>\rho_{c}(\beta)$, signaling that finite clusters cannot carry a density greater than $\rho_{c}(\beta)$.
Now whether $A_{\rho,\beta}(\bar{\mu}(\rho,\beta))$ is finite or not depends on the asymptotic form of $q_{n}$. If
@@
$A_{\rho,\beta}(\mu)=e^{\beta\mu-\rho\hat{v}(0)}+\sum_{n=3}^{\infty}\alpha_{n}e^{-n\beta\delta_{n}(\beta)}e^{n[\beta\mu-\beta\epsilon(\beta)-\rho\hat{v}(0)]}\qquad(q_{1}=1,q_{2}=0).$ (14)
Because $n\delta_{n}(\beta)=o(n)$, $\beta\mu-\beta\epsilon(\beta)-\rho\hat{v}(0)=0$ determines the convergence radius via
@@
independent of $\rho$! We see that the fate of infinite clusters depends on the correction $e^{o(n)}$ to the exponential decay or growth of $|q_{n}|$; the whole difficulty of proving the transition lies in this fact. In this respect, the appearance of infinite clusters is analogous to the appearance of infinite permutation cycles in systems of interacting bosons [18].
It cannot be ruled out that for certain interactions $|q_{n}|\sim\exp\{-c_{1}n^{1+c_{2}}\}$ where $c_{1},\,c_{2}>0$. If this holds true, the convergence radius of the series $A_{\rho,\beta}(\mu)$ is infinite, there are only finite clusters, and there are no classical phase transitions at all densities. This may be the case for long-range repulsive forces which, as previously noted [18], can also prevent the occurrence of infinite permutation cycles and thus of Bose-Einstein condensation.
One general characteristic remains to be mentioned. Phase diagrams, whether experimental or theoretical, exhibit a critical point $(\rho^{*}=\rho_{c}(\beta^{*}),\beta^{*})$ on the $(\rho,\beta)$ plane, such that for $\beta<\beta^{*}$ the gas does not condense and only a single fluid phase exists. In our study this translates into the fact that the function $\bar{A}_{\beta}$ tends towards infinity as $\beta$ approaches a certain value $\beta^{*}>0$ from above. For $\beta<\beta^{*}$, Eq. (12) can be solved with $\mu<\bar{\mu}(\rho,\beta)$, where $\mu-\bar{\mu}(\rho,\beta)<0$ approaches zero as $\rho$ increases.